# A cautionary tale of model misspecification and identifiability

### v2

@@ 1 Introduction
While resolving identifiability and potentially improving the precision of parameter estimates, model simplification can exacerbate model misspecification [11]. To visualise the effect of model misspecification on parameter inference, consider a process characterised by the generalised logistic growth model for cell density $u(t)$
@@
where f(u^)f($\hat${u}) is the crowding function and $K$ is the carrying capacity. The specific form of f(u^)f($\hat${u}) in Eq. 1 is commonly referred to as Richards model [12]. In our previous work [4], we demonstrate that calibration of the full parameter set $(r,K,\beta)$ to data can lead to issues with practical identifiability, manifesting as a strong dependence between the exponent parameter $\beta$ and the low-density growth rate parameter $r$. It is, however, common practice to assume that density-dependent growth processes follow (at least approximately) the canonical logistic growth model, corresponding to Eq. 1 with fixed $\beta=1$. Although not always the result of an identifiability analysis, the resulting simplified model yields improved parameter identifiability and estimate precision, however the introduced misspecification is potentially problematic for parameter estimation accuracy [13].
@@
Issues relating to misspecification can be detected from residual analysis [4, 15] or, should a sufficient quantity of data be available, through model validation. Fundamentally, however, issues relating to misspecification can arise from the inability of a (potentially) simple and identifiable model to capture the true underlying dynamics [16]. Even relaxing assumptions on $\beta$ and calibrating Richards model to cell growth data is potentially problematic: both in terms of the prior that must be placed on the purely phenomenological exponent $\beta$ and the assumed parametric form for the crowding function. Significant effort has focused on relaxing assumptions related to the functional form of terms in differential equations [17]. Symbolic regression, a parametric approach, attempts to “learn” the functional form of an ordinary or partial differential equation using a library of potential terms, applying sparse regression to avoid overfitting [18]. Non-parametric approaches, meanwhile, include those that represent one or more terms using a Gaussian process or a neural network [19, 20, 21, 22, 23, 24, 25], or through expansion or spline representation of any unknown functions [26, 27, 28, 29]. Recently introduced in the context of differential equation modelling in biology, Biological Informed Neural Networks (BINNs) [30] constrain neural networks through a biologically informed model structure; for example, a reaction-diffusion equation with density-dependent diffusivity and reaction rate. All the aforementioned approaches alleviate model misspecification; however, they are often heavily parameterised and data intensive. With a modest amount of experimental data, it remains unclear how these functional-form-free approaches affect the identifiability of key quantities of interest.
@@ 4 Discussion and conclusions
Complex biological systems necessitate correspondingly complex models. Available data, however, are often limited: even relatively simple mathematical models can be practically non-identifiable. A persistent challenge in mathematical biology is to balance model complexity and identifiability without exacerbating issues related to model misspecification. Complex or over-parameterised models may provide superior fits, but the lack of parameter identifiability limits biological insight. Simpler misspecified models, on the other hand, may lead to precise but biased parameter estimates. The primary goal, therefore, should not be the precise estimation of model parameters, but the accurate quantification of the uncertainty in parameter estimates.

### v3

@@ 1 Introduction
While resolving identifiability and potentially improving the precision of parameter estimates, model simplification can exacerbate model misspecification [11]. To visualise the effect of model misspecification on parameter inference, we first consider a process characterised by the generalised logistic growth model for cell density $u(t)$
@@
where f(u^)f($\hat${u}) is the crowding function and $K$ is the carrying capacity. The specific form of f(u^)f($\hat${u}) in Eq. 1 is commonly referred to as Richards model [12]. In our previous work [4], we demonstrate that calibration of the full parameter set $(r,K,\beta)$ to data can lead to issues with practical identifiability, manifesting as a strong dependence between the exponent parameter $\beta$ and the low-density growth rate parameter $r$. The specific choice of $\beta$ is often based on convention, and not data analysis: in cell and cancer biology, the canonical logistic model ($\beta=1$) and to a lesser extent the Gompertz model $(\beta\rightarrow 0)$ are routine [13, 14, 15]. The von Bertalanffy model is historically associated with fish growth, although selection between various parametric models is now accepted practice in this field [16]. Although not always the result of an identifiability analysis, restricting study to a simplified (e.g. logistic) model can yield improved parameter identifiability and estimate precision, however the introduced misspecification is potentially problematic for parameter estimation accuracy [17].
@@
Issues relating to misspecification can be detected from residual analysis [4, 19] or, should a sufficient quantity of data be available, through model validation. Fundamentally, however, issues relating to misspecification can arise from the inability of a (potentially) simple and identifiable model to capture the true underlying dynamics [20]. Even relaxing assumptions on $\beta$ and calibrating Richards model to cell growth data is potentially problematic: both in terms of the prior that must be placed on the purely phenomenological exponent $\beta$ and the assumed parametric form for the crowding function. Significant effort in recent years has focused on relaxing assumptions related to the functional form of terms in differential equations [21]. Symbolic regression, a parametric approach, attempts to “learn” the functional form of an ordinary or partial differential equation using a library of potential terms, applying sparse regression to avoid overfitting [22]. Non-parametric approaches, meanwhile, include those that represent one or more terms using a Gaussian process or a neural network [23, 24, 25, 26, 27, 28, 29], or through expansion or spline representation of any unknown functions [30, 31, 32, 33]. Recently introduced in the context of differential equation modelling in biology, Biological Informed Neural Networks (BINNs) [34] constrain neural networks through a biologically informed model structure; for example, a reaction-diffusion equation with density-dependent diffusivity and reaction rate. All the aforementioned approaches alleviate model misspecification; however, they are often heavily parameterised and data intensive. With a modest amount of experimental data, it remains unclear how these functional-form-free approaches affect the identifiability of key quantities of interest.
@@ 4 Discussion and conclusions
Complex biological systems necessitate correspondingly complex models. Available data, however, are often limited: even relatively simple mathematical models can be practically non-identifiable. A persistent challenge in mathematical biology is to balance model complexity and identifiability without exacerbating issues related to model misspecification. Complex or over-parameterised models may provide superior fits, but the lack of parameter identifiability limits biological insight. Simpler misspecified models, on the other hand, may lead to precise but biased parameter estimates and potentially misleading conclusions. This phenomenon is analogous to Simpson’s paradox [52] in statistics: incorporating additional freedom in a misspecified model can, potentially substantially, alter estimates of other parameters. The primary goal, therefore, should not be the precise estimation of model parameters, but the accurate quantification of the uncertainty in parameter estimates.